# Environment-Driven Online LiDAR-Camera Extrinsic Calibration

### v2

@@ IX More Visualization Results
In this section, we provide more detailed visualization results of EdO-LCEC through LiDAR and camera data fusion on MIAS-LCEC datasets. Fig. 10 demonstrates that the LiDAR point cloud and camera image aligned almost perfectly along the textural and geometric edges, proving the accuracy of our calibration results. It can also be observed that our method performs well in challenging conditions, particularly under poor illumination and adverse weather, or when few geometric features are detectable. This impressive data fusion result further validates the critical role of LiDAR-camera extrinsic calibration. The seamless integration of spatial and textural modalities equips robots with a human-like perceptual ability to interpret and interact with their surroundings. Our environment-driven approach bridges the gap toward the flexible and adaptive machines often envisioned in science fiction.

### v3

@@ IX More Visualization Results
In this section, we present detailed visualization results of EdO-LCEC through LiDAR-camera data fusion on the KITTI Odometry and MIAS-LCEC datasets. An example of the fusion result using our calibrated extrinsic parameters for the left camera in KITTI odometry is illustrated in Fig. 11. The LiDAR point clouds and camera RGB images are precisely aligned under the estimated calibration parameters, producing highly consistent geometric and photometric correspondences. As shown in Fig. 12, the LiDAR projections and camera images exhibit near-perfect alignment along both textural and structural edges, confirming the sub-pixel accuracy achieved by our method. Furthermore, our algorithm maintains high robustness under challenging conditions, including poor illumination, adverse weather, and scenes with sparse or ambiguous geometry. The ability to preserve geometric consistency across such conditions highlights the strong generalization capacity of our environment-driven strategy. Beyond quantitative metrics, these visualization results intuitively demonstrate the practical impact of EdO-LCEC. By enabling seamless fusion of 3D spatial geometry and 2D visual texture, it allows robots to perceive, localize, and interact with their surroundings in a manner closer to human visual understanding. This perceptual synergy between LiDAR and camera further paves the way toward adaptive, context-aware robotic systems capable of reliable operation in complex real-world environments.